I can't remember how to do this...
Solve. 15x^4-14x^2+3=0

Answer by stanbon(75887)

Solve. 15x^4-14x^2+3=0
-----------------
15x^4 - 9x^2-5x^2 + 3 = 0
3x^2(5x^2-3) - (5x^2 - 3) = 0
---
(5x^2-3)(3x^2-1) = 0
---------------
5x^2 = 3 or 3x^2 = 1
x^2 = 3/5 or x^2 = 1/3
----
x = +-sqrt(15)/5 or x = +-sqrt(3)/3
=======================================
Cheers,
Stan H.
=======================================